When 279855 is added to a number, it became 152051 less than the greatest number of six digits. Find the number?

Grade:6

1 Answers

Vaibhav Singh
7 years ago
The greatest six digit number is 999999.Let the required number be x. Then, 279855 + x = 999999 - 152051.Therefore, x = 568093. Hence the required number is 568093.